A movie star, unwilling to give his age, posed the following riddle to a gossip columnist: "6 years ago, I was 11 times as old as my daughter. Now I am 5 times as old as she is." How old is the movie star?
Found 2 solutions by ikleyn, ankor@dixie-net.com:
Answer by ikleyn(52776)

.
A movie star, unwilling to give his age, posed the following riddle
to a gossip columnist: "6 years ago, I was 11 times as old as my daughter.
Now I am 5 times as old as she is." How old is the movie star?
~~~~~~~~~~~~~~~~~~

Let x be the daughter's age now.

Then the movie star's age now is 5x, according to the problem.


6 years ago, the star was (5x-6) years old, 
while her daughter was x-6 years old.


Write an equation for their ages 6 years ago

    5x-6 = 11*(x-6).


Simplify and find x

    5x - 6 = 11x - 66

    66 - 6 = 11x - 5x

       60  =    6x

        x  =    60/6 = 10.


Thus, the daughter's age now is 10 years; hence, the movie star is 5*10 = 50 years old now.


ANSWER.  The movie star is 50 years old now.

Solved.



Answer by ankor@dixie-net.com(22740)

A movie star, unwilling to give his age, posed the following riddle to a gossip columnist:
"6 years ago, I was 11 times as old as my daughter.
Now I am 5 times as old as she is."
How old is the movie star?
:
let a = his present age
let d = daughter's present age
Write an equation for each statement
"
"6 years ago, I was 11 times as old as my daughter."
a - 6 = 11(d-6)
a - 6 = 11d - 66
a = 11d - 66 + 6
a = 11d - 60
and
" Now I am 5 times as old as she is."
a = 5d
replace a with 5d in the 1st equation
5d = 11d - 60
5d - 11d = -60
-6d = - 60
d = -60/-6
d = +10 is the daughters age
then
a = 5(10)
a = 50 yrs is his age
:
:
Check solution in the 1st statement
50 - 6 = 11(10-6)
44 = 11*(4)